OpenStudy (anonymous):

A cylinder has a radius of 10 cm and a height of 8 cm. What is the volume of the cylinder? I got 2513.27 but it's wrong

OpenStudy (anonymous):

volume is base times height the base is the area of the circle. What is formula for area of circle?

OpenStudy (anonymous):

A= pie times r^2

OpenStudy (anonymous):

correct, no plug in 10 for the radius

OpenStudy (anonymous):

now*

OpenStudy (anonymous):

31.4

OpenStudy (anonymous):

alright then 31.4 * 8

OpenStudy (anonymous):

25.12

OpenStudy (anonymous):

wait.... it is 100 *3.14 which is 314.159

OpenStudy (anonymous):

sorry, so it is 314.159*8

OpenStudy (anonymous):

So you got it right

OpenStudy (anonymous):

where did you get the 100?

OpenStudy (anonymous):

r^2 = 10^2 = 100

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

I don't know why that is not right. Maybe check decimal places?

OpenStudy (anonymous):

also remember units

OpenStudy (anonymous):

So its 2513?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thats not one of the options on the test was the problem because I got that same answer

OpenStudy (anonymous):

do you remember any of the choices?

OpenStudy (anonymous):

Yeah ill put in a picture of the questions and chocies

OpenStudy (anonymous):

ohhhhhhh

OpenStudy (anonymous):

They want your answers in terms of pi

OpenStudy (anonymous):

so just divide our answer by 3.14159

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

800!!

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

no problem